# Excitonic Theory of the Ultrafast Optical Response of 2D-Quantum-Confined Semiconductors at Elevated Densities

### v1

@@ I Introduction
Rabi oscillations Rabi (1937), while first having been observed in atomic two-level systems Knight and Milonni (1980), later in quantum dots Danckwerts et al. (2006); Förstner et al. (2003) or assisted by optical cavities Dominici et al. (2014), also occur in confined semiconductors such as quantum wells without an optical cavity Schülzgen et al. (1999); Giessen et al. (1998); Cundiff et al. (1994); Kuhn and Rossi (1992); Binder et al. (1990). In these structures, however, Rabi oscillations are not only determined by the usual Pauli blocking but also by Coulomb interactions between optically excited electrons and holes. Rabi oscillations in quantum wells have been successfully predicted by the semiconductor Bloch equations in Hartree-Fock limit (SBE) and observed in experiments Lindberg and Koch (1988); Schülzgen et al. (1999); Binder et al. (1999). However, if – in materials exhibiting strong electron-hole interactions – higher correlations beyond the Hartree-Fock limit and incoherent interaction with the phonon bath need to be taken into account, the numerical effort of the SBEs quickly increases Steinhoff et al. (2025). Also, the Hartree-Fock-limit lacks any possibility to resolve correlated exciton-exciton effects Takayama et al. (2002); Buck et al. (2004); Steinhoff et al. (2018); Katsch et al. (2019, 2020); Tang and Ning (2025) or an accurate description of intraband processes Dignam et al. (2014); Parks et al. (2013); Axt et al. (1996a). Therefore, for material parameters, where Coulomb interaction dominates over the light-matter interaction, such as atomically thin transition metal dichalcogenide monolayers, an excitonic theory, i.e., a formulation in electron-hole pair operators Katsch et al. (2018); Axt and Stahl (1994); Victor et al. (1995), is better suited.
@@
In the past, several excitonic theories valid in the regime of elevated densities have been proposed: By using a momentum-independent formulation using the Usui transformation Usui (1960); Rochat et al. (2000); Wang and Dignam (2009); Wang et al. (2007), by using a localized Wannier basis Axt et al. (2001a), by employing a strict coherent limit Östreich and Knorr (1994); Axt et al. (2001b); Katsch et al. (2020, 2019) or via a bosonic non-equilibrium Green’s function method focusing on the dynamics of optically inactive excitons via nonlinear light-matter interaction Yang et al. (2004). Other approaches regarding exciton dynamics in optical cavities introduce additional Hamiltonians to account for fermionic Pauli-blocking in an otherwise bosonic excitonic description Takemura et al. (2015, 2016); Combescot et al. (2007a) or stay in a strict coherent limit Rose et al. (2025); Denning et al. (2022a, b).
@@
Hence, a comprehensive and microscopic momentum-resolved theory for Wannier-Mott excitons Wannier (1937) developing a consistent description of simultaneous Coulomb, light-matter and exciton-phonon interaction in an undoped quantum-confined semiconductor, which is able to bridge the gap between the coherent Takayama et al. (2002); Buck et al. (2004); Katsch et al. (2019, 2020); Deckert et al. (2025) and incoherent regime Thränhardt et al. (2000); Selig et al. (2018); Brem et al. (2020); Katzer et al. (2023); Policht et al. (2023) in the low-density limit as well as in the elevated-density regime beyond the third-order susceptibility Katzer et al. (2024, 2023); Erkensten et al. (2021); Steinhoff et al. (2021), is still missing.
@@
In this manuscript, we develop an excitonic approach for Wannier excitons up to fourth-order dynamics-controlled truncation (DCT) Axt and Stahl (1994); Victor et al. (1995) for up to three-exciton correlations needed to describe the optical response from the coherent to the incoherent regime at elevated exciton densities.
@@ II Hamiltonians
The total Hamiltonian $\hat{H}$ reads Lindberg and Koch (1988); Haug and Koch (2009); Jahnke et al. (1997):
@@
where $E_{v/c,\mathbf{k}}^{\xi}$ $E_{v/c,\mathbf{k}}$ is the valence/conduction band dispersion in effective-mass approximation $E_{v/c,\mathbf{k}}^{\xi}=E_{v/c}^{\xi}\mp\frac{\hbar^{2}\mathbf{k}^{2}}{2m_{h/e}^{\xi}}$ with valence/conduction band edge $E_{v/c}^{\xi}$ and hole/electron masses $m_{h/e}^{\xi}$ at momentum $\mathbf{k}$ and spin-valley index $\xi$. $v^{(\dagger)}$$c^{(\dagger)}$ are the valence/conduction band annihilation (creation) operators. The Coulomb Hamiltonian Jahnke et al. (1997) consists of an electron-hole contribution $\hat{H}_{\text{Coul-eh}}$:
@@
Here, $V_{\mathbf{q}}=\frac{e^{2}}{\mathcal{A}}\int\mathrm{d}z\,\mathrm{d}z^{\prime}\,\left|\xi(z)\right|^{2}G_{\mathbf{q}}\left(z,z^{\prime}\right)\left|\xi(z^{\prime})\right|^{2}$ is the screened quantum-confined Coulomb potential, where $e$ is the elementary charge, $\mathcal{A}$ is the sample area, $\xi(z)$ is the confinement wave function and $G_{\mathbf{q}}(z,z^{\prime})$ is the Green’s function solving the generalized Poisson equation of the corresponding sample geometry including the dielectric environment, cf. Eq. (S68) in the Supplementary Material (SM). Coulomb exchange interaction – not to be confused with the usual contributions due to fermionic exchange – Qiu et al. (2015); Yu and Wu (2014); Combescot et al. (2023), which causes corrections to the exciton Qiu et al. (2015) or biexciton dispersion Kwong et al. (2021), intervalley scattering Selig et al. (2019, 2020) and the biexciton fine structure Steinhoff et al. (2018); Torche and Bester (2021), as well as Auger scattering Steinhoff et al. (2021); Erkensten et al. (2021) and Dexter interaction Dogadov et al. (2025); Berghäuser et al. (2018); Bernal-Villamil et al. (2018), are neglected, since we focus on the most dominant (direct) Coulomb interaction processes needed to explain recent experiments Schäfer et al. (2025). However, these interaction processes can be straightforwardly included by adding the corresponding Hamiltonians.
@@
Here, $\mathbf{E}_{\mathbf{Q}}=\frac{1}{\mathcal{A}}\int\mathrm{d}z\,|\xi(z)|^{2}\mathbf{E}_{\mathbf{Q}}(z)$ is the quantum-confined optical field, where $\mathbf{E}_{\mathbf{Q}}(z)$ solves the wave equation in momentum space of the corresponding sample geometry Knorr et al. (1996) and $\mathbf{d}^{cv}$ is the transition dipole moment in low-wavenumber approximation Aversa and Sipe (1995); Gu et al. (2013). As we excite excitonic states just below the free-particle band gap, we neglect any intraband contributions Huttner et al. (2017). Note, that intraband light-matter interaction can induce light-driven transport of optically excited electron-hole pairs Oliaei Motlagh et al. (2021), intraexcitonic transitions Rice et al. (2013); Stroucken et al. (2021) or dynamical localization Meier et al. (2000, 1995), even throughout the total Brillouin zone if the exciting pulse is strong enough Langer et al. (2018), constituting the key ingredient for lightwave electronics Borsch et al. (2023). However, in such a regime, the effective-mass approximation and, hence, a projection on excitonic states solving the Wannier equation, cf. Eq. (13), is not possible anymore, since we would have to take into account the full connected band structure throughout the total Brillouin zone.
@@ III Excitonic Equations of Motion
The excitonic equations of motion are derived as follows. We start by deriving Heisenberg’s equations of motion $\mathrm{i}\hbar\partial_{t}\hat{A}=[\hat{A},\hat{H}$] for the electron-hole pair operators $\hat{A}$ of interest, which can contain an arbitrary but equal number of conduction/valence band annihilation (creation) operators $c^{(\dagger)}$$v^{(\dagger)}$. However, after evaluating the commutator $\hat{A},\hat{H}$], the appearing operator products on the right-hand side of the corresponding equation of motion do not necessarily contain an equal number of electron and hole operators. Therefore, we project all operator products, which contain an unequal number of electron and hole operators onto electron-hole pairs via the unit-operator method Katsch et al. (2018); Ivanov and Haug (1993). E.g., electron and hole occupations are expanded according to (using a compound index $i=\{\mathbf{k}_{i},\xi_{i}\}$):
@@
Note, that the unit-operator method corresponds exactly to the contraction theorem in Ref. Victor et al. (1995). Then, we evaluate the expectation values via the correlation expansion technique Fricke (1996); Kira and Koch (2006). Within this framework, the expectation values of the operator products are expanded as follows ($\hat{A}_{i}$ is an arbitrary fermionic creation or annihilation operator):
@@
Here, the subscript ”c” denotes the correlated part of the expectation value on the corresponding operator level, which cannot be factorized in lower contributions anymore. The sign of each term is determined by the amount of commutations necessary to disentangle the operator products. The correlated expectation values behave antisymmetric with respect to pairwise exchange:
@@
We formulate our theory always in the truly correlated quantities $\langle\hat{A}\rangle_{\text{c}}$, which are obtained by subtracting all lower contributions from the total expectation value Fricke (1996); Hoyer et al. (2002):
@@
This definition provides the clearest physical picture, since it ensures, that the corresponding quantum kinetics only include fully correlated contributions on the corresponding level (singlet, doublet, triplet etc.) and no lower-order correlations mix in. In particular, it ensures a clear separation of coherent and incoherent excitonic effects. To determine the relevant microscopic quantities, we perform a dynamics-controlled truncation (DCT) Axt and Stahl (1994); Victor et al. (1995). This approach systematically organizes the quantum-mechanical hierarchy problem in powers of the optical field $\mathbf{E}$. Within the Heisenberg-equations-of-motion method, Eq. (7) provides the possibility to introduce the dynamical order $m$ of the expectation value of an excitonic operator $\hat{A}_{n}$ of order $n$:
$\langle\hat{A}^{(n)}\rangle=\mathcal{O}(\mathbf{E}^{m}),$ (11)
@@
where $m\geq n$. Here, $n$ is the total number of valence or conduction band creation and annihilation operators of $\hat{A}^{(n)}$. Therefore, if one wants to calculate the optical response of a semiconductor, one first has to fix the desired dynamical order $m$ within the optical field and subsequently collects all occurring (correlated) expectation values of order $n\leq m$. To determine the order $m$ of a specific expectation value, one counts all occurring electron-hole pairs and optical fields on the right-hand side of the corresponding Heisenberg equations of motion. After we set up all equations motion in the electron-hole picture, we perform an expansion in excitonic wave functions afterwards. This is crucial, since approaches, which apply an excitonic picture already on the Hamiltonian level are prone to miss important corrections due to the fermionic substructure of the excitons Combescot et al. (2007b) within higher orders of the optical field.
We note, that the DCT-approach is similar but not equal to a strict perturbation expansion in powers of the optical field Aversa and Sipe (1995), which is often done for calculating, e.g., high-harmonic responses, as the DCT-approach always delivers a closed set of equations, while the strict perturbation expansion does not.
@@
Figure 1: Scheme of excitonic transitions in Eq. (13) at the $K$ and $K^{\prime}$ valley (left) and excitonic occupations from Eq. (15) in intravalley configuration at the $K$ and $K^{\prime}$ valleys (right) in a monolayer TMDC.
@@
Throughout this work, we restrict the excitonic occupations/intraexcitonic coherences to carrying total momenta of $2\mathbf{Q}$, since we assume spatially homogeneous optical excitation.
As introduced by Moskalenko Moskalenko (1958) and Lampert Lampert (1958) and later explicitly calculated by Ivanov using a variational approach Ivanov et al. (1978), the two-exciton coherence of order $n=2$ are defined as Katsch et al. (2019); Schumacher et al. (2006); Takayama et al. (2002); Schäfer et al. (1996), cf. Fig. 2(left):
@@
where $\Phi^{*}{\vphantom{\varphi}}_{B,\pm,\zeta,\mu,\nu,\mathbf{Q}}^{\text{L},\xi_{1},\xi_{2},\xi_{3},\xi_{4}}$ is the two-exciton wave function solving the two-exciton Schrödinger equation with quantum number $\zeta$ and configuration $\pm$, cf. Eq. (S18). $(S_{\pm})_{\mu,\nu,\mathbf{Q},\rho,\eta,\mathbf{Q}^{\prime}}^{\xi_{1},\xi_{2},\xi_{3},\xi_{4}}(\mathbf{0})$ is the projection matrix element arising due to the expansion in symmetric (++) and antisymmetric (-) configurations with respect to electron exchange, cf. Eq. (S23). The subscript $\xi/\xi^{\prime}$” with respect to the electron configuration in the excitonic wave functions denotes an averaging over the electron configurations $\xi$ and $\xi^{\prime}$, if they are different, i.e., the corresponding excitonic wave functions obey the averaged Wannier equation, cf. Eq. (S1). This averaging procedure is necessary to expand the two-exciton correlations in excitonic wave functions. Throughout this work, we use the term ”two-exciton coherence” to generally denote all bound and unbound two-exciton states with $\zeta=b$ or $\zeta\neq b$, while we use the term ”two-exciton continuum”, if we specifically designate the continuum states with $\zeta\neq b$, and we use the term ”bound biexciton”, if we specifically designate the bound two-exciton state $\zeta=b$. Note, since we only consider optically excited excitonic transitions at zero center-of-mass momentum $\mathbf{Q}=\mathbf{0}$, the two-exciton states are restricted to a total momentum of zero, which greatly reduces the relevant two-exciton phase space.
@@
First rigorously calculated by Axt Axt and Stahl (1994), the exciton-two-exciton transitions of order $n=3$ are defined as, cf. Fig. 2(right) and Fig. 3:
@@
where $\Phi^{*}{\vphantom{\varphi}}_{Z,\pm,\zeta,\mu,\nu,\rho,\mathbf{Q}^{\prime}}^{\text{L},\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}(\mathbf{Q})$ is the exciton-two-exciton wave function solving the exciton-two-exciton Schrödinger equation with quantum number $\zeta$ and configuration $\pm$, cf. Eq. (S19). We use the term ”exciton-two-exciton transition” to generally denote all bound and unbound exciton-two-exciton states, while we use the term ”exciton-two-exciton continuum”, if we specifically designate the continuum states, and we use the term ”bound exciton-biexciton”, if we specifically designate the bound exciton-two-exciton state. Note, that the exciton-two-exciton transition is a different many-body correlation than a triexciton, as described in Refs. Meier et al. (2003); Bartels et al. (1995), which is at least relevant in a fifth-order response. Also note, that the main difference to the two-exciton transitions $B$ in Eq. (18) is the additional center-of-mass-$\mathbf{Q}$-dependence, i.e., the exciton-two-exciton transition carries a total momentum of $2\mathbf{Q}$ (excitonic center-of-mass momentum), as their source is of partly incoherent nature. Also note, that there is some ambiguity in the interpretation of Eq. (20), since $Z\sim\langle\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\rangle_{\text{c}}$ can be equally interpreted as a correlated coherence between an excitonic transition and a two-exciton transition $\langle(\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}})(\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})\rangle_{\text{c}}$ or as an occupation-assisted excitonic transition $\langle(\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})(\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})\rangle_{\text{c}}$. However, both of these contributions are taken into account.
@@
Throughout this work, we disregard correlated eight-operator expectation values of order $n=4$ (quadruplets), which are expected to induce excitonic incoherent Coulomb scattering in the sense of exciton-exciton collisions. While it has been shown, that Coulomb scattering in the electron-hole picture beyond the Hartree-Fock limit Bányai et al. (1998) significantly suppresses Rabi oscillations in GaAs QWs, they can be recovered if the electron and hole occupation dynamics are formulated in a dressed-state basis Ciuti et al. (2000). This is the reason, why the semiconductor Bloch equations applied to material systems with weak Coulomb interaction already in Hartree-Fock limit successfully reproduced actual measured Rabi oscillations Schülzgen et al. (1999); Cundiff et al. (1994). Therefore, the neglection of incoherent Coulomb scattering in the study of Rabi oscillations in confined semiconductors such as conventional GaAs-like QWs, where light-matter interaction is of similar strength compared to Coulomb interaction, is justified as a first approximation. Future studies might shed more light on this issue.

### v2

@@ I Introduction
Rabi oscillations [97], while first having been observed in atomic two-level systems [66], later in quantum dots [30, 42] or assisted by optical cavities [36], also occur in confined semiconductors such as quantum wells without an optical cavity [107, 44, 29, 70, 18]. In these structures, however, Rabi oscillations are not only determined by the usual Pauli blocking but also by Coulomb interactions between optically excited electrons and holes. Rabi oscillations in quantum wells have been successfully predicted by the semiconductor Bloch equations in Hartree-Fock limit (SBE) and observed in experiments [80, 107, 19]. However, if – in materials exhibiting strong electron-hole interactions – higher correlations beyond the Hartree-Fock limit and incoherent interaction with the phonon bath need to be taken into account, the numerical effort of the SBEs quickly increases [117]. Also, the Hartree-Fock-limit lacks any possibility to resolve correlated exciton-exciton effects [119, 23, 115, 59, 60, 122] or an accurate description of intraband processes [34, 93, 7]. Therefore, for material parameters, where Coulomb interaction dominates over the light-matter interaction, such as atomically thin transition metal dichalcogenide monolayers, an excitonic theory, i.e., a formulation in electron-hole pair operators [58, 6, 127], is better suited.
@@
In the past, several excitonic theories valid in the regime of elevated densities have been proposed: By using a momentum-independent formulation using the Usui transformation [126, 100, 129, 130], by using a localized Wannier basis [8], by employing a strict coherent limit [92, 9, 60, 59] or via a bosonic non-equilibrium Green’s function method focusing on the dynamics of optically inactive excitons via nonlinear light-matter interaction [133]. Other approaches regarding exciton dynamics in optical cavities introduce additional Hamiltonians to account for fermionic Pauli-blocking in an otherwise bosonic excitonic description [121, 120, 27] or stay in a strict coherent limit [102, 32, 33].
@@
Hence, a comprehensive and microscopic momentum-resolved theory for Wannier-Mott excitons [132] developing a consistent description of simultaneous Coulomb, light-matter and exciton-phonon interaction in an undoped quantum-confined semiconductor, which is able to bridge the gap between the coherent [119, 23, 59, 60, 31] and incoherent regime [124, 110, 22, 62, 95] in the low-density limit as well as in the elevated-density regime beyond the third-order susceptibility [61, 62, 39, 116], is still missing.
@@
In this manuscript, we develop an excitonic approach for Wannier excitons in a few-band effective-mass model at the symmetry points of the Brillouin zone up to fourth-order dynamics-controlled truncation (DCT) [6, 127] for up to three-exciton correlations needed to describe the optical response from the coherent to the incoherent regime at elevated exciton densities.
@@ II Hamiltonians
The total Hamiltonian $\hat{H}$ reads [80, 47, 55]:
@@
where $E_{v/c,\mathbf{k}}^{\xi}$ $E_{v/c,\mathbf{k}}$ is the valence/conduction band dispersion in effective-mass approximation $E_{v/c,\mathbf{k}}^{\xi}=E_{v/c}^{\xi}\mp\frac{\hbar^{2}\mathbf{k}^{2}}{2m_{h/e}^{\xi}}$ with valence/conduction band edge $E_{v/c}^{\xi}$ and hole/electron masses $m_{h/e}^{\xi}$ at momentum $\mathbf{k}$ and spin-valley index $\xi$. $v^{(\dagger)}$$c^{(\dagger)}$ are the valence/conduction band annihilation (creation) operators. The Coulomb Hamiltonian [55] consists of an electron-hole contribution $\hat{H}_{\text{Coul-eh}}$:
@@
Here, $V_{\mathbf{q}}=\frac{e^{2}}{\mathcal{A}}\int\mathrm{d}z\,\mathrm{d}z^{\prime}\,\left|\zeta(z)\right|^{2}G_{\mathbf{q}}\left(z,z^{\prime}\right)\left|\zeta(z^{\prime})\right|^{2}$ is the screened quantum-confined Coulomb potential, where $e$ is the elementary charge, $\mathcal{A}$ is the sample area, $\zeta(z)$ is the confinement wave function and $G_{\mathbf{q}}(z,z^{\prime})$ is the Green’s function solving the generalized Poisson equation of the corresponding sample geometry including the dielectric environment, cf. Eq. (S68) in the Supplementary Material (SM). Coulomb electron-hole exchange interaction – not to be confused with the usual contributions due to fermionic exchange – [96, 135, 28], which causes corrections to the exciton [96] or biexciton dispersion [72], intervalley scattering [112, 111] and the biexciton fine structure [115, 125], as well as Auger scattering [116, 39] and Dexter interaction [35, 15, 16], are neglected, since we focus on the most dominant (direct electron-hole, electron-electron and hole-hole) Coulomb interaction processes needed to explain recent experiments [103]. However, these interaction processes can be straightforwardly included by adding the corresponding Hamiltonians.
@@
Here, $\mathbf{E}_{\mathbf{Q}}=\frac{1}{\mathcal{A}}\int\mathrm{d}z\,|\zeta(z)|^{2}\mathbf{E}_{\mathbf{Q}}(z)$ is the quantum-confined optical field, where $\mathbf{E}_{\mathbf{Q}}(z)$ solves the wave equation in momentum space of the corresponding sample geometry [67] and $\mathbf{d}^{cv}$ is the transition dipole moment in low-wavenumber approximation [4, 46]. As we excite excitonic states just below the free-particle band gap, we neglect any intraband contributions [52]. Note, that intraband light-matter interaction can induce light-driven transport of optically excited electron-hole pairs [91], intraexcitonic transitions [99, 118] or dynamical localization [87, 84], even throughout the total Brillouin zone if the exciting pulse is strong enough [76], constituting the key ingredient for lightwave electronics [21]. However, in such a regime, the effective-mass approximation and, hence, a projection on excitonic states solving the Wannier equation, cf. Eq. (14), is not possible anymore, since we would have to take into account the full connected band structure throughout the total Brillouin zone.
We note, that we disregard any exciton-phonon scattering causing incoherent scattering and thermalization [124, 110, 22, 62, 61], since the goal of this work is to study the impact of exciton-exciton interactions via Coulomb interaction in its pure form.
@@ III Excitonic Equations of Motion
The excitonic equations of motion are derived as follows. We start by deriving Heisenberg’s equations of motion $\mathrm{i}\hbar\partial_{t}\hat{A}=[\hat{A},\hat{H}$] for the electron-hole pair operators $\hat{A}$ of interest, which can contain an arbitrary but equal number of conduction/valence band annihilation (creation) operators $c^{(\dagger)}$$v^{(\dagger)}$. However, after evaluating the commutator $\hat{A},\hat{H}$], the appearing operator products on the right-hand side of the corresponding equation of motion do not necessarily contain an equal number of electron and hole operators. Therefore, we project all operator products, which contain an unequal number of electron and hole operators onto electron-hole pairs via the unit-operator method [58, 53]. E.g., electron and hole occupations are expanded according to (using a compound index $i=\{\mathbf{k}_{i},\xi_{i}\}$):
@@
Note, that the unit-operator method corresponds exactly to the contraction theorem in Ref. [127]. Then, we evaluate the expectation values via the correlation expansion technique [43, 65]. Within this framework, the expectation values of the operator products are expanded as follows ($\hat{A}_{i}$ is an arbitrary fermionic creation or annihilation operator):
@@
Here, the subscript “c” denotes the correlated part of the expectation value on the corresponding operator level, which cannot be factorized in lower contributions anymore. The sign of each term is determined by the amount of commutations necessary to disentangle the operator products. The correlated expectation values behave antisymmetric with respect to pairwise exchange:
@@
We formulate our theory always in the truly correlated quantities $\langle\hat{A}\rangle_{\text{c}}$, which are obtained by subtracting all lower contributions from the total expectation value [43, 51]:
@@
This definition provides the clearest physical picture, since it ensures, that the corresponding quantum kinetics only include fully correlated contributions on the corresponding level (singlet, doublet, triplet etc.). In particular, it ensures a clear separation of coherent and incoherent excitonic effects. To determine the relevant microscopic quantities, we perform a dynamics-controlled truncation (DCT) [6, 127]. The DCT is a method to systematically truncate the quantum-mechanical hierarchy problem, where a specific correlation order is linked to a certain order of the optical field $\mathbf{E}$. Notably, the Heisenberg-equations-of-motion method, Eq. (7), provides the possibility to classify the dynamics of an occurring expectation value of an arbitrary operator $\hat{A}^{(n)}$ containing a total number $n$ of valence or conduction band creation and annihilation operators (whichever number is higher) in powers of the optical field $m$:
$\mathrm{i}\hbar\partial_{t}\langle\hat{A}^{(n)}\rangle=\mathcal{O}(\mathbf{E}^{m}),$ (11)
where $m\geq n$. More specifically, $m$ is the dynamical field order of a given equation of motion of $\hat{A}^{(n)}$ and $n$ is a measure of the electron-hole pair order of $\hat{A}^{(n)}$. $m$ and $n$ are not necessarily equal. Therefore, if the overall optical response of a semiconductor up to a certain order $m$ in a perturbation approach of nonlinear optics within the optical field is required, one first has to fix the desired dynamical order $m$ and subsequently collects all occurring (correlated) expectation values of order $n\leq m$. To determine the order $m$ of the dynamical equation of a specific expectation value, one counts all occurring electron-hole pairs and optical fields on the right-hand side of the corresponding Heisenberg equations of motion.
For example, if we are interested in the second-order optical response, we fix $m=2$ and collect all correlated expectation values of order $n\leq 2$, which is the excitonic transition $P=\langle\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\rangle_{\text{c}}$ with $n=1$, the excitonic occupation $N=\langle\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\rangle_{\text{c}}$ with $n=2$ and the two-exciton transition $B=\langle\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\rangle_{\text{c}}$ with $n=2$, where we use a short-hand notation without indices: The full expressions can be found in Eq. (14), Eq. (16) and Eq. (19), respectively. Then, we calculate the corresponding equations of motion of $P$, $B$ and $N$, which we truncate by using the condition $m=2$ on the right-hand side. Schematically, the equations of motion read:
$\begin{split}\mathrm{i}\hbar\partial_{t}P=&\,EP-\hbar\Omega,\\ \mathrm{i}\hbar\partial_{t}N=&\,0,\\ \mathrm{i}\hbar\partial_{t}B=&\,E_{B}B+VP^{2},\end{split}$ (12)
where $E$ is the excitonic energy, $\Omega$ is the Rabi frequency induced by the exciting optical field $\mathbf{E}$, $E_{B}$ is the two-exciton energy and $V$ is the Coulomb-induced exciton-exciton interaction. Again, the corresponding full versions of Eq. (12) can be found in Eq. (22), Eq. (28) and Eq. (30). Since the optical response is fully characterized by the dynamics of the excitonic transitions $P$ and higher-order correlations $N$ and $B$ do not couple back to the excitonic transitions $P$ in the first line in Eq. (12), the optical response in second order is identical to the first order.
@@
After we set up all equations of motion in the electron-hole picture, we perform an expansion in excitonic wave functions afterwards. This is crucial, since approaches, which apply an excitonic picture already on the Hamiltonian level are prone to miss important corrections due to the fermionic substructure of the excitons [26] within higher orders of the optical field.
@@
Figure 1: Scheme of excitonic transitions in Eq. (14) at the $K$ and $K^{\prime}$ valley (left) and excitonic occupations from Eq. (16) in intravalley configuration at the $K$ and $K^{\prime}$ valleys (right) in a TMDC ML.
@@
Throughout this work, we restrict the excitonic occupations/intraexcitonic coherences to carry total momenta of zero, since we assume spatially homogeneous optical excitation, where no net momentum is injected into the excitonic system. Note, that the total momentum of $N\vphantom{N}_{\mu,\nu,\mathbf{Q}}^{\xi,\xi^{\prime}}$ is not equal to the center-of-mass momentum of an electron-hole pair $\mathbf{Q}$ but rather corresponds to the center-of-mass momentum of a pair of two excitons, which is zero.
As introduced by Moskalenko [90] and Lampert [74] and later explicitly calculated by Ivanov using a variational approach [54], the two-exciton transition of order $n=2$ are defined as [59, 108, 119, 104], cf. Fig. 2(left):
@@
where $\Phi^{*}{\vphantom{\varphi}}_{B,\pm,\zeta,\mu,\nu,\mathbf{Q}}^{\text{L},\xi_{1},\xi_{2},\xi_{3},\xi_{4}}$ is the two-exciton wave function solving the two-exciton Schrödinger equation with quantum number $\zeta$ and configuration $\pm$, cf. Eq. (S18). $(S_{\pm})_{\mu,\nu,\mathbf{Q},\rho,\eta,\mathbf{Q}^{\prime}}^{\xi_{1},\xi_{2},\xi_{3},\xi_{4}}(\mathbf{0})$ is the projection matrix element arising due to the expansion in symmetric (++) and antisymmetric (-) configurations with respect to electron exchange, cf. Eq. (S23). The subscript $\xi/\xi^{\prime}$” with respect to the electron configuration in the excitonic wave functions denotes an averaging over the electron configurations $\xi$ and $\xi^{\prime}$, if they are different, i.e., the corresponding excitonic wave functions obey the averaged Wannier equation, cf. Eq. (S1). This averaging procedure is necessary to expand the two-exciton correlations in excitonic wave functions. Throughout this work, we use the term “two-exciton transition” to generally denote all bound and unbound two-exciton states with $\zeta=b$ or $\zeta\neq b$, while we use the term “two-exciton continuum”, if we specifically designate the continuum states with $\zeta\neq b$, and we use the term “bound biexciton”, if we specifically designate the bound two-exciton state $\zeta=b$. Note, since we only consider optically excited excitonic transitions at zero center-of-mass momentum $\mathbf{Q}=\mathbf{0}$, the two-exciton states are restricted to a total momentum of zero, which greatly reduces the relevant two-exciton phase space.
@@
First rigorously calculated by Axt [6], the exciton-two-exciton transitions of order $n=3$ are defined as [8, 20, 85, 13], cf. Fig. 2(right) and Fig. 3:
@@
where $\Phi^{*}{\vphantom{\varphi}}_{Z,\pm,\zeta,\mu,\nu,\rho,\mathbf{Q}^{\prime}}^{\text{L},\xi_{1},\xi_{2},\xi_{3},\xi_{4},\xi_{5},\xi_{6}}(\mathbf{Q})$ is the exciton-two-exciton wave function solving the exciton-two-exciton Schrödinger equation with quantum number $\zeta$ and configuration $\pm$, cf. Eq. (S19). We use the term “exciton-two-exciton transition” to generally denote all bound and unbound exciton-two-exciton states, while we use the term “exciton-two-exciton continuum”, if we specifically designate the continuum states, and we use the term “bound exciton-biexciton”, if we specifically designate the bound exciton-two-exciton state. Note, that the exciton-two-exciton transition is a different many-body correlation than a triexciton, as described in Refs. [86, 12], which is at least relevant in a fifth-order response. Also note, that the main difference to the two-exciton transitions $B$ in Eq. (19) is the additional center-of-mass-$\mathbf{Q}$-dependence, as their source is of partly incoherent nature. Also note, that there is some ambiguity in the interpretation of Eq. (21), since $Z\sim\langle\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\rangle_{\text{c}}$ can be equally interpreted as a correlated coherence between an excitonic transition and a two-exciton transition $\langle(\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}})(\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})\rangle_{\text{c}}$ or as an occupation-assisted excitonic transition $\langle(\hat{c}^{\dagger}\hat{v}^{\vphantom{\dagger}}\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})(\hat{v}^{\dagger}\hat{c}^{\vphantom{\dagger}})\rangle_{\text{c}}$. However, both of these contributions are taken into account.
@@
Throughout this work, we disregard correlated eight-operator expectation values of the form $\langle\hat{c}^{\dagger}_{1}\hat{v}_{2}^{\vphantom{\dagger}}\hat{c}^{\dagger}_{3}\hat{v}_{4}^{\vphantom{\dagger}}\hat{v}^{\dagger}_{5}\hat{c}_{6}^{\vphantom{\dagger}}\hat{v}^{\dagger}_{7}\hat{c}_{8}^{\vphantom{\dagger}}\rangle_{\text{c}}$ of order $n=4$ (quadruplets), which describe two-exciton occupations or density-density correlations giving rise to excitonic incoherent Coulomb scattering in the sense of thermalization-causing exciton-exciton collisions [106, 105]. While it has been shown, that Coulomb scattering in the electron-hole picture beyond the Hartree-Fock limit [11] significantly suppresses Rabi oscillations in GaAs QWs, they can be recovered if the electron and hole occupation dynamics are formulated in a dressed-state basis [25]. This is the reason, why the semiconductor Bloch equations applied to material systems with weak Coulomb interaction already in Hartree-Fock limit successfully reproduced actual measured Rabi oscillations [107, 29]. Therefore, the neglection of incoherent Coulomb scattering in the study of Rabi oscillations in confined semiconductors such as conventional GaAs-like QWs, where light-matter interaction is of similar strength compared to Coulomb interaction, is justified as a first approximation. Future studies might shed more light on this issue.